A sample of a gas is in a sealed container. The pressure of the gas is 475 torr , and the temperature is 12 ∘C . If the temperature changes to 84 ∘C with no change in volume or amount of gas, what is the new pressure, P2, of the gas inside the container

Using the same sample of gas (P1 = 475 torr , T1 = 12 ∘C ), we wish to change the pressure to 4750 torr with no accompanying change in volume or amount of gas. What temperature T2, in Celsius, is needed to reach this pressure?
Answer #1

1)

Given:

Pi = 475 torr

Ti = 12.0 oC

= (12.0+273) K

= 285 K

Tf = 84.0 oC

= (84.0+273) K

= 357 K

use:

Pi/Ti = Pf/Tf

475 torr / 285.0 K = Pf / 357.0 K

Pf = 595 torr

Answer: 595 torr

2)

Given:

Ti = 12.0 oC

= (12.0+273) K

= 285 K

Pi = 475 torr

Pf = 4750 torr

use:

Pi/Ti = Pf/Tf

475.0 torr / 285.0 K = 4750.0 torr / Tf

Tf = 2850 K

= (2850 - 273) oC

= 2577 oC

Answer: 2.58*10^3 oC